OpenStudy (anonymous):

log 5(4x-13) = log 5 (x+2) solve this

OpenStudy (oaktree):

You know that 4x-13 has to = x+2, so just solve.

OpenStudy (oaktree):

3x=15, and x=5.

OpenStudy (anonymous):

duh. -__-

OpenStudy (anonymous):

@Williamskates are you okie here?

OpenStudy (lgbasallote):

do you still not get it @Williamskates ? since you bumped it..

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

a property \[\LARGE \log_a b = \log_a c \Longleftrightarrow b = c\] that just means that when two logarithms with two different powers but the same base..the owers are equal...since both your logarithms have base 5...then you just equate the powers to one another arriving with \[4x - 13 = x + 2\]